Assume the Black-Scholes framework for options pricing. You are a portfolio manager and already have a long position in AppleAssume the Black-Scholes framework for options pricing. You are a portfolio manager and already have a long position in AppleO 100% O 50% 0 0% 0 15.4%
Answer #1

As per Black Scholes Model, value of put option

p = K*e^(-r*t) * N(-d2) - S *N(-d1)

where

Spot price S =$200

Strike price K = $180.15

risk free rate r =0.08

standard deviation s=0.3

time in years t= 1

d1= ( ln(S/K) + (r + s^2/2) *t ) / (s*t^0.5) = 0.765

d2 =d1-s*t^0.5 = 0.465

So, N(-d1) = 0.2221 (Can be calculated by NORMSDIST function in EXCEL)

and N(-d2) =0.3209

Putting the above values , we get the value of put option as

p = $8.949

As , the value of put option with strike $180.15 is $8.949

the call option with strike price $180.15 cant have a premium of $8.949 as it has an intrinsic value of around $20 and also time value for one year

So, possible strike prices are $200, $270.15 and $280.15

The price of a call option as per BSM model

C=S*N(d1)-K*e^(-r*t) * N(d2), putting the values in this equation for K =$200 we get

C = $31.42

The price of a call option as per BSM model

C=S*N(d1)-K*e^(-r*t) * N(d2), putting the values in this equation

For K =$200 we get

C = $31.42

For K =$270.15 we get

C = $8.951

So, the correct option is $270.15